# Safe MPC Alignment with Human Directional Feedback

### v1

@@ III-C1 Representation of Human Intended Safety
We consider the scenario of human-robot coexistence, where a human is supervising the robot for a given task. The user has his/her intent of robot safety. In Safe MPC (1), we hypothesize that the user intended safety can be represented by the set
@@
Here, $\boldsymbol{\theta}_{H}$ is user implicit intent parameter picked from an implicit set $\Theta_{H}\subset\mathbb{R}^{r}$, i.e., $\boldsymbol{\theta}_{H}\in\Theta_{H}$. $\Theta_{H}$ is implicit to the human user and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot behavior to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve the safety of the robot.
We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
III-C2 Human Online Directional Correction
We consider a human user who gives correction to the robot in an online fashion while the robot is running with the most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, denote $i$ to be the index of learning update, and suppose the robot policy currently is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. At the time steps $k_{i}$ of rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, a human user gives directional correction $\boldsymbol{a}_{k_{i}}$ in robot action space at its state $\boldsymbol{x}_{k_{i}}$. After receiving $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) and obtain the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$. Then, the robot continues with the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ until the next user possible correction. This process of robot MPC policy update and user correction repeats until the convergence of the policy update or the user is satisfied with the robot motion. An illustration of one update is shown in Fig. 2.
@@ III-C2 Human Online Directional Correction
In the above-stated human feedback process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later in Section IV-A). For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go. Also, we assume that the user feedback is applied in the action space. For some robotic systems, this assumption can be readily met, such as turning the driving wheels for autonomous vehicles. In other cases, the correction can be applied through a dedicated (hardware or algorithmic) interface to the action space; such as contact forces from end-effector to joint torque, and teleoperation device for remote control. In those cases, $\boldsymbol{a}$ is the correction signal after the interface. The reason why we only consider the correction in action space is that otherwise the corrections in robot state space can lead to dynamics unfeasibility or uncontrollability for under-actuated robotic systems [7, 51].
@@ III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned policy update procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}$. However, it is possible that the robot may still exit the user-intended safety zone during the correction process. In such cases, a human user may trigger an emergency stop to stop the robot motion and potentially restart it from a desired safety region. This emergency stop mechanism is common in real-world robotic applications. Until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless. For instance, correcting an autonomous vehicle after it has been crashed is futile.
@@ III-C4 Parametric Safety Constraints
In the safe MPC (1), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$. We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex safety region, which can be alternatively represented by the intersection of multiple constraints (e.g., a polytope safety set defined by multiple hyperplanes). In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the extensive prior work on learning cost/reward [4, 7, 52, 53, 54]. Specifically, we parameterize $\boldsymbol{g}_{\theta}$
${g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})=\phi_{0}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\theta_{1}\phi_{1}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\\ +\theta_{2}\phi_{2}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1}),$ (5)
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, the $\boldsymbol{\theta}$ is scaling ambiguous, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions. We do not impose any constraints on the choice of feature vector as long as it is differentiable. For example, $\boldsymbol{\phi}$ can be neural features. Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
@@ III-C5 Problem of Safe MPC Alignment
Following the procedure of online human correction and robot policy update described in Subsection III-C and given a weighted-feature safety constraint (5), the interest of this work is to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ with each human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., the safety constraint in the robot safe MPC policy will be aligned with human intent.
IV Algorithm Overview for Safe MPC Alignment
IV-A Hypothesis on User Directional Feedback
In this session, we will introduce a hypothesis on user directional correction, which lays the foundation for our alignment framework proposed subsequently. Informally, we have the following intuitions on user feedback. First, user directional corrections, on average, improve the robot safety. The similar assumption is the basis for reinforcement learning with human feedback [55]. Second, user correction on robot safety is typically proactive and typically happens before the robot enters into the true unsafe region. In other words, human correction is made when the robot operates within $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a ground vehicle after its collision with obstacles. Third, with the second intuition, user feedback is typically safety-margin aware: the closer a robot gets to the boundary of the unsafe region, the more significant user safety correction will be. This is observed in our prior user study [7]. In the following, we will formalize those intuitions.
Suppose that robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is solved from (3) with parameter $\boldsymbol{\theta}$, and a human user applies directional correction $\boldsymbol{a}$ at first time step 00. Define
@@ IV-A Hypothesis on User Directional Feedback
Formally, we assume there is a nonzero-volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit intent set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the user directional correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. Concretely, it means $\mathbf{a}$ on averages points in a direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ evaluated at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
@@
Here, ${\operatorname{\mathbb{E}}}_{\mathbf{a}}(\mathbf{a})=\mathbf{\bar{a}}$ and the equality is due to the second term $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ is not dependent on $\mathbf{a}$. $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ denotes the full gradient of $B$ with respect to $\boldsymbol{u}_{0:T-1}$ evaluated at trajectory $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$. Similar notations apply to $J$ and $g$ below. Since state sequence $\boldsymbol{x}_{0:T}$ is the rollout of $\boldsymbol{u}_{0:T-1}$ on dynamics, the differentiation goes through the dynamics equation.
One of the important observations to (7) is that the magnitude of the user correction $\mathbf{a}$ does not affect the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot safeMPC plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the user directional correction always accounts for half of the action space, as shown in Figure 3. Extending (7) leads to
@@
(I) The first and second terms on the left side of (8) correspond to the user’s correction toward lowering the cost and improving the safety value of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. $\gamma$ serves as a weight between two improvements.
(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) indicates that user’s correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more significant a human user correction will be to drive the system back to safety interior. This property is consistent with our previous intuition.
(III) The above hypothesis (7) requires that $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place, i.e.,
@@
The requirement corresponds to the third intuition — a user typically makes correction only when the robot is within user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session. Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning pipeline developed in the next session.
@@ IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
To find an implicit $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that at any learning iteration $i=0,1,2,....$, the robot monitors and updates a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. At initial, the robot may choose a hypothesis space $\Theta_{0}$ that is large enough to contain $\bar{\Theta}_{H}$. Specifically, one can choose
$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{h}\}\subset\mathbb{R}^{r},$ (10)
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$ are the given bounding vectors. By specifying a $\Theta_{0}$ large enough, we assume
@@
Later, we will that our alignment algorithm can self-access the failure of learning in case of misspecification, where (11) is not met.
Recall the procedure of human correction in Section III-C and Fig. 2. At $i$-th learning update, $i=1,2,3...$, the robot safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$. By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypothesis (7) and (9)
@@
which we refer to as the cutting set. Using $\mathcal{C}_{i}$, one can update the hypothesis space
@@
The above update can be viewed as a step of using the cutting set (12) to cut the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible. The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used in the next iteration. With the above update, we have the following result.
@@
As shown by Fig. 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction. Therefore, we name our alignment method Hypothesis Space Cutting. As a summary, we present the three main steps of the hypothesis space cutting updates below.
@@
At $i$th iteration of the Hypothesis Space Cutting, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction is used to induce the cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,...$.
From (13) and by following the above hypothesis space cutting method, one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq\Theta_{2}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq\operatorname{{Vol}}\Theta_{2}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing along with user corrections. It is hoped that the volume of search space is strictly decreasing, such that new $\boldsymbol{\theta}_{i+1}$ has a very good chance of falling inside $\bar{\Theta}_{H}$, which means a successful safety alignment between the robot and human. We will later (in Section V) theoretically show that the above Safe MPC Alignment method, within finite interactions $K$, will either

### v2

@@ III-C1 Representation of Human Intended Safety
Consider a human-robot coexistence task, where a human is supervising the robot motion. The user has his/her intent of robot motion safety. In Safe MPC (1), we hypothesize user-intended safety can be represented by a subset of the trajectory space:
@@
Here, $\boldsymbol{\theta}_{H}$ is intent parameter from an implicit set $\boldsymbol{\theta}_{H}\in\Theta_{H}\subset\mathbb{R}^{r}$. $\Theta_{H}$ is implicit to human and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve its safety. We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
III-C2 Human Directional Correction
We consider human gives correction to a robot in an online fashion while the robot is executing its most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. During rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a sparse human directional correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
Here, we do not assume precise timing $k_{i}$ or duration of the human correction due to the sparsity or intermittency of human feedback. In other words, human can apply one or multiple corrections at any time in robot rollout time steps for any duration as he/she finds necessary. But we here assume human directional corrections $\boldsymbol{a}_{k_{i}}$ are in robot action space. For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with steering wheel and pedals. In some other applications, the action space correction can be achieved via specific mapping or interface. For example, to correct a torque-controlled robot arm, human’s correction on robot end-effector can be mapped into joint torque via Jacobian transpose matrices. The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system [20].
After receiving each $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its MPC policy. Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user corrections. This process repeats until the convergence of the policy update or the user is satisfied with the robot motion. An illustration of one update is shown in Fig. 2.
@@ III-C2 Human Directional Correction
In the above learning process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later). For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go. As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
@@ III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned learning procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$. However, it is possible that the robot may still exit the user-intended safety region. In such cases, a human user may trigger an emergency stop mechanism to stop the robot motion and potentially restart it from a desired safety region. In this sense, all trajectories before the human triggers the emergency stop are considered safe in human intent. In other words, until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless. For instance, correcting a vehicle after it is crashed is futile.
@@ III-C4 Parametric Safety Constraints
In the safe MPC (4), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$. We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex (non-convex) safety region, which can be alternatively represented by the intersection of multiple constraints. In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the prior work on learning cost/reward [4, 20, 50, 51, 52]. Specifically, we parameterize $\boldsymbol{g}_{\theta}$
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}(\boldsymbol{\xi})+\theta_{1}\phi_{1}(\boldsymbol{\xi})+\theta_{2}\phi_{2}(\boldsymbol{\xi})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{\xi}),$ (5)
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions. The dimension (number) $r$ of features can be chosen such that $\boldsymbol{\phi}$ has enough expressiveness to represent the human-intended safety constraint. We do not impose any constraints on the choice of feature vector as long as it is differentiable, for example, $\boldsymbol{\phi}$ can be neural features. In practice the feature vector can be chosen by trial and error, i.e., starting from a few features and gradually adding more until the learned constraint is satisfactory to the human user. Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
@@ III-C5 Problem of Safe MPC Alignment
Following the procedure of human correction and robot policy update in Subsection III-C and given a weighted-feature safety constraint (5), we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
IV Overview of Safe MPC Alignment Algorithm
IV-A Hypothesis on Human Directional Corrections
Informally, we have the following intuitions on human feedback. First, human corrections, on average, improve the robot safety. Second, human correction on robot safety is typically proactive and happens before the robot leaves the human-intended safety set. In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a vehicle after its crash. Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task objective. For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target. In the following, we will formalize those intuitions.
Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space. Define
@@ IV-A Hypothesis on Human Directional Corrections
We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
@@
[…]
(7) indicates that the magnitude of $\mathbf{a}$ does not affect the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the allowable human corrections satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig. 3. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$, by extension of the gradient, (7) becomes
@@
(I) The first and second terms on the left side of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively. $\gamma$ serves as a weight between two improvements.
(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human user emphasize on improving the robot’s safety than minimizing the objective. This property is consistent with our previous intuition.
(III) (7) implies $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place:
@@
This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session. Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning next.
@@ IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that for each human directional correction received $i=0,1,2,....$, the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. At initial, the robot may choose a large enough hypothesis space $\Theta_{0}$. Specifically, one can choose $\Theta_{0}$ as a box space
$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{u}\}\subset\mathbb{R}^{r},$ (10)
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, resepctively, and the inequalities applied element-wise. By saying a $\Theta_{0}$ large enough, we assume
@@
Later, we will show that our method can self-access the learning failure in case of misspecification, where (11) is not met.
Recall the procedure of human correction in Section III-C and Fig. 2. At $i$-th learning iteration, $i=1,2,3...$, the robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$. By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypotheses (7) and (9):
@@
which we refer to as the cutting set. Note that the first inequality $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is deduced by multiplying -1 on both sides of (7). Using $\mathcal{C}_{i}$, one can update the hypothesis space
@@
The above update can be viewed as a step of using (12) to cut (intersect with) the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible. The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used for the next iteration. With the above update, we have the following result.
@@
As shown in Fig. 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction. Therefore, we name our alignment method Hypothesis Space Cutting. As a summary, we present the three main steps of our algorithm below.
@@
At $i$th learning iteration, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,...$. From (13), one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing after each user correction. It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of falling inside $\bar{\Theta}_{H}$ — a successful safety alignment between the robot and human.
For the above Safe MPC Alignment method, we will later (in Section V) show that within finite number $K$ of corrections, the method will either